suppose a customer has income of N$120 per period and faces prices, price of X = 2 and price of Z =3. Her goal is to maximize her utility, described by the functional U = 10X0.5Z0.5..What is the consumers budget constraint?.. State the formula for finding marginal utilities for goods X and Z.. Calculate the marginal utilities for goods X and Z.. State utility maximizing condition for the consumer... Calculate the utility maximizing bundle (X*, Z*)


1
Expert's answer
2022-03-15T12:09:36-0400

A) The consumer's budget constraint is given as:

"2X+3Z=120"


B) The marginal utility of X, "MU_X" is given as:

"MU_X=\frac{dU}{dX}"

The marginal utility of Z, "MU_Z" is given as:

"MU_Z=\frac{dU}{dZ}"


C) Given that

"u=10X^{0.5}Z^{0.5}"

"MU_X=5X^{-0.5}Z^{0.5}"

"MU_Z=5X^{0.5}Z^{-0.5}"


D) The utility maximizing condition is given as:

"\frac{MU_X}{MU_Z}=\frac{P_X}{P_Z}"


E) "u=10X^{0.5}Z^{0.5}" subject to the constraint

"120-2X-3Z"

"\frac{5X^{-0.5}Z^{0.5}}{5X^{0.5}Z^{-0.5}}=\frac{2}{3}"

"\frac{Z}{X}=\frac{2}{3}"

"X=1.5Z"

Substituting x Into the constraint:."120-2(1.5Z)-3Z=0"

"120=6Z"

"Z^*=20"

"X^*=1.5(20)=30"